OpenStudy (anonymous):

Erika was working on solving the exponential equation 50^x = 17; however, she is not quite sure where to start. Using complete sentences, describe to Erika how to solve this equation and how solving would be different if the bases were equal.

OpenStudy (anonymous):

any help?

OpenStudy (kirbykirby):

Um.. I can write the mathematical steps, then write what you do ..

OpenStudy (anonymous):

This whole lesson has lost me I'm usally pretty decent with math but I really dont get this at all

OpenStudy (kirbykirby):

\(50^x=17\) \(\log 50^x=\log 17\) \(x \log 50 = \log 17\) \(x=\large\frac{\log 17}{\log 50}\)

OpenStudy (anonymous):

and if the bases where equal like 4=8 2^2=2^3?

OpenStudy (kirbykirby):

you would just have to compare the exponents. Say you had \(8^x=4\), then you can write it as \((2^3)^x=2^2\), or \(2^{3x}=2^2\) In this case, you just need to compare the exponents and solve \(3x=2\)

OpenStudy (kirbykirby):

Although technically you can still you use logs.

OpenStudy (anonymous):

I get it now well more than I did before lol thank you so much!

OpenStudy (kirbykirby):

:) Cool! JUst remember, when solving an exponential equation, just log both sides because you can solve for x easily that way.

OpenStudy (anonymous):

ok thank you

OpenStudy (kirbykirby):

np